The sum of three consecutive integers is 72. Find the integers.



1
Expert's answer
2022-01-18T10:32:54-0500

Solution:

Let n - 1 - first consecutive integer

Let n - second consecutive integer

Let n + 1 - third consecutive integer

So, the sum of three consecutive numbers:

(n - 1) + n + (n + 1) = 72

n - 1 + n + n + 1 = 72 

3n = 72

n = 24

n - 1 = 23 

n + 1 = 25 


Check:

(n - 1) + n + (n + 1) = 72

23 + 24 + 25 = 72

72 = 72


Answer:

The three consecutive integers whose sum is 72 are 23, 24, and 25